# Gamma-ray Bursts as Distance Indicators by a Machine Learning Approach

### v1

@@ 4 Summary and Conclusion
We tested an ensemble with several ML algorithms to determine the redshift of GRBs based on characteristics related to their prompt and, for the first time, the plateau emission. Some afterglow characteristics (duration of the plateau emission, its flux, and the decay after the plateau) are among the most powerful predictors. A successful application of these methods to the Swift existing GRB catalog can nearly double the sample of GRBs with observed redshift and observed plateau, thus yielding a much larger sample. This, in turn, has allowed us to determine the LF and the density rate evolution of GRBs with unprecedented accuracy. Using a sample of 147 GRBs from the Swift satellite with known redshifts, we demonstrate the SuperLearner’s capabilities to provide accurate $z_{pred}$ that agree with the $z_{obs}$ with a Pearson correlation coefficient of $r=0.93$, and MSE=$0.21$. This result makes GRBs powerful distance estimators, improving upon other existing methods. Indeed, we obtained an increase of $61\%$ in the Pearson correlation of the $z_{pred}$ and $z_{pred}$. Instead, previous works, which omitted plateau parameters and used only one ML method, achieved a correlation of $r=0.57$. Results from applying this method to a large number of LGRBs have cast light on the controversy on the discrepancy between the density rate of LGRBs and the cosmological SFR and to accurately measure the GRB LFs. Our results mirror similar deviations in the formation rates of LGRBs found by Tsvetkova et al. (2021); Lloyd-Ronning et al. (2019); Yu et al. (2015); Petrosian et al. (2015); Pescalli et al. (2016). Notably, Pescalli et al. (2016) suggests a deviation of 2.6 sigma when considering the star formation history of Madau & Dickinson (2014). Furthermore, Dainotti et al. (2021) also arrived at a similar conclusion regarding SGRBs at low redshifts, even though the delayed SFR has been accounted for. Our primary conclusion, which aligns with findings reached by Petrosian & Dainotti (2023), is that the population of LGRBs at low redshift includes a component of compact merger progenitors, similar to the SGRBs. A different explanation for the excess of GRBs at low redshift, originally proposed by Yoon et al. (2006), is related to the increase in metallicity with cosmic time, which prevents a larger fraction of massive stars from producing GRBs.
@@ 5 Acknowledgement
We are particular grateful to Michal Ostrowski for his precious discussion during the data analysis and results, to Trevor Nelson for his preliminary analysis during the summer internship program at Stanford University supported by SLAC, to Simona Scimemi for the preliminary analysis related only to prompt emission variables during the winter internship 2018 at INAF Bologna, to Monica Brora on the very preliminary analysis on random forest. We also grateful to Macjei Bilicki for the useful discussion on the parameter space division. We thank particularly Douglas Figberg for his interesting discussion on selection biases related to this procedure and the way on how to solve them. M.G.D is particularly grateful for the support received as RIKEN pionieer research during her stay in Riken. M.G.D is grateful to the Gill Ferrandes and Donald Warren for their interesting discussions on the method of SuperLearner. We are also grateful to Trevor Hastie for useful suggestions to avoid overfitting and piece of the code for the 10-fold cross-validation. We are also grateful for the initial discussion on the project to Dmitry Maleshev. M.G.D. acknowledges the funding from MINIATURA2, Numer: 2018/02/X/ST9/03673. M.G.D also acknowledges Prof. Peter Behroozi for the insightful discussion on how to avoid overfitting in our machine learning models. This research was also support by the Visibility and Mobility module of the Jagiellonian University (Grant number: WSPR.WSDNSP.2.5.2022.5) and the NAWA STER Mobility Grant (Number: PPI/STE/2020/1/00029/U/00001). We acknowledge funding from Swift GI Cycle XIX, grant number 22-SWIFT22-0032. This work was also supported by the Polish National Science Centre grant UMO-2018/30/M/ST9/00757 and by Polish Ministry of Science and Higher Education grant DIR/WK/2018/12.
@@ A.1 Machine learning
We have used predictive data mining, also called supervised learning. It is based on prior knowledge of a ‘training’ data set on which we can build models to predict the relations in a new “test” data set. Examples are regression tools, such decision trees, random forests, and gradient boosting. These allow us to estimate complicated non-linear relationships between the response and explanatory variables as well as high-order interactions between predictors. Such non-parametric methods are very powerful when the data set contains many observations. However, they suffer from the so-called “curse of dimensionality”, which limits the number of explanatory variables one can efficiently estimate for a small sample size. As a result, the fully non-parametric machine learning methods allow using only a limited number of predictors, e.g., when estimating redshifts based on small GRB training sets. On the other hand, when important explanatory variables are identified by other methods or expert knowledge, these methods allow the construction of efficient predictors without assumptions on the underlying form of the relevant correlations among the variables. In our analysis, we have used supervised ML tools that employ regression. Within regression methods, we have used parametric, semi-parametric, and non-parametric approaches. We give a brief description of each model below.
A.1.1 Feature selection
@@
The GRB sample contains a total of nine predictors. Four belong to the prompt: $Fluence$, $T_{90}$, $\log(Peakflux)$, and $Photon$ $Index$. The rest belong to the plateau and afterglow: $T_{a}$, $F_{a}$, $\alpha$, $\beta$, and $NH$. To pick out the best predictors of redshift among these 9 variables, we applied the Least Absolute Shrinkage and Selection Operator (LASSO) feature selection technique. The LASSO technique applies a shrinkage method for linear regression by forcing the $\ell^{1}$ norm (sum of the magnitude of all vectors in the given space) of the result vector to be $\leq\lambda$, where $\lambda$ is a positive number known as the tuning parameter. This form of penalization enables the model to choose specific features while nullifying others by assigning their coefficients a value of 0 (Tibshirani, 1996). The tuning parameter determines the degree of shrinkage applied to the estimated vector. Consequently, when using fewer features, the model becomes more interpretable and generally exhibits lower prediction errors than the full model. For our analysis, we use the GLMNET function with the LASSO selection feature (Hastie et al., 2017; Tibshirani et al., 2012). We pick the $\lambda.1se$ value, the maximum $\lambda$ value for which the error is within one standard deviation (Friedman et al., 2010a), and its corresponding coefficients for the features. Thus, we drop $T_{90}$ and $Fluence$ while keeping the rest of the 7 variables. The coefficients assigned by LASSO to each of them are displayed in Fig. 6, and we choose only the non-zero coefficient features. The LASSO selected features are also presented in Fig. 1.
@@
Figure 6: LASSO-based feature selection, showing the weights assigned to the different variables.
@@ A.1.2 The Generalized Additive Model
In the semi-parametric generalized additive models (GAM), $Y$ is related to the predictors, $x_{1},\ldots,x_{p}$ through the additive model $f(x_{1},\ldots,x_{p})=\beta_{0}+\sum_{i=1}^{p}{\beta_{i}f_{i}(x_{i})}$. The functions $f$ may have some specified parametric form (e.g., a polynomial) or be estimated non--parametrically, simply as ‘smooth functions.’ In this method, interactions between different predictors can be included. An interaction may arise when there are two or more explanatory variables, and the influence of some of them on the response variable depends on the other explanatory variables. In practice, interactions make it more difficult to predict the consequences of changing the value of a variable. If any smooth functions were allowed in model fitting, then the maximum likelihood estimation of such models would result in over-fitting. Thus, the models are usually fit by penalized likelihood maximization, in which the model is modified by adding a penalized likelihood estimation, namely a penalty for each smooth function, penalizing its irregularities.
We use GAM to determine an analytical formula that best estimates the $\log(z+1)$ value. It should be noted that we present results in log scale because of the log scale transformations we applied to the majority of the variables and because the best model is indeed trained in $\log(1+z)$. In addition, Z = $z+1$ is a more natural parametrization of the cosmological evolution. The parameters of this formula are the LASSO selected predictors and their cross multiplicative terms (terms defined as the products among the 7 chosen variables). A total of 16,511 formulas were generated considering all possible combinations of all seven explanatory variables and the product of the pairs of these variables.
@@
$\log(z+1)=(\log(Peak)^{2}+\log(NH)^{2}+PI^{2}+\beta^{2}+\log(Peak)+\alpha)^{2}+\log(F_{a})+\log(T_{a})+\beta+PI+\log(NH)+\log(F_{a})^{2}+\log(T_{a})^{2}+\alpha^{2}$ (A1)
$\log(z+1)=(\log(Peak)^{2}+\log(NH)^{2}+\beta^{2}+\log(Peak)+\alpha)^{2}+\log(F_{a})+\log(T_{a})+\beta+PI+\log(NH)+\log(F_{a})^{2}+\log(T_{a})^{2}+\alpha^{2}+PI^{2}$ (A2)
$\log(z+1)=(\log(Peak)^{2}+\log(NH)^{2}+\beta^{2}+\log(Peak)+\log(NH))^{2}+\log(F_{a})+\log(T_{a})+\alpha+\beta+PI+\log(F_{a})^{2}+\log(T_{a})^{2}+\alpha^{2}+PI^{2}$ (A3)
@@ A.1.3 Step-AIC
Step-AIC is a parametric approach that chooses algorithms based on the Akaike Information Criterion (AIC) Sakamoto et al. (1986); Venables et al. (2002), which, given a collection of models, estimates the quality of each model relative to each other. The StepAIC model works by fitting a formula to which it adds (or subtracts) explanatory variables iteratively while minimizing the AIC.
@@ A.1.4 The Generalized Linear Model
Generalized Linear Model (GLM) Nelder & Wedderburn (1972) is a parametric method that estimates the mean of the response variable’s distribution using a linear model of the explanatory variables. The response variable is connected to the desired target variable (GRB redshift) via a link function. This is an extension of the ordinary linear model. Using the appropriate link functions, we can consider non-Gaussian target variable distributions, such as Poisson or Gamma functions. GLM adopts the iteratively weighted least squares method for maximum likelihood estimation to fit these functions. Here, we assign all three formulas (Eq. A1, A2, and A3) as the desired fitting functions, while the link function is Gaussian.
@@ A.2 Outlier removal with M-estimator
To remove outliers that might not represent the full sample, we apply the Robust Linear regression Method (RLM) with the M-estimator. This is an alternative technique to the ordinary least squares method, which for a univariate response variable, $Y$ and a vector $X=(X_{1},\ldots,X_{p})$ of explanatory variables $X$, attempts to fit the following relation: $f(x_{1},\ldots,x_{p})=\beta_{0}+\sum_{i=1}^{p}{\beta_{i}x_{i}}$. The ordinary least squares method attempts to minimize the square of the residuals (called the $L_{2}$ norm regression). However, this gives outliers of the data set a higher weight, which significantly affects the results of the regression fit. In contrast, the M-estimator attempts to minimize the sum of a function of the residuals. The function chosen for our analysis is the Huber Function (Huber, 1964).
@@
Iterated re-weighted least squares do the M-estimator fitting, and the weights, ranging from 0 to 1, are assigned based on the residuals (Huber, 1996). RLM is used for the detection of highly influential observations. We are using the implementation of RLM as described in the MASS package of R (Venables & Ripley, 2002).
@@
Here, we fit our dataset with Eq. A1 using the M-estimator and choose a cut-off weight of 0.5, eliminating GRBs with a weight $\leq 0.5$. The threshold of 0.5 was chosen to ensure we remove less than 5% of our sample as outliers. With the M-estimator, the sample size is reduced from 150 to 147.
A.3 The SuperLearner and the Relative importance of the features
@@
SuperLearner is an algorithm that uses internal 10-fold cross-validation (10fCv) to estimate the performance of multiple ML models or the same model with different settings. It minimizes the estimated risk (a measure of model accuracy) by minimizing the RMSE. To test the performance of our results and estimate the prediction error, besides the 10fCV, we also used an external layer of cross-validation called nested cross-validation. Namely, we performed 10fCV repeatedly $100$ times, with the folds of the cross-validation step being randomized for each iteration. This type of cross-validation involves partitioning the sample into complementary subsets (in this case, $10$), fitting the model on 9 out of $10$ folds (training sets) of the full data set, and validating the analysis on the remaining 1 testing set. The procedure is repeated iteratively so that each set of the $10$ will be used as a test set and the remaining as training sets, and the prediction results are averaged over the number of runs (in this case $100$). This step allows us to remove the randomness inherent to the cross-validation procedure and stabilize the results. Then, SuperLearner creates an optimal weighted average of those models, e.g., an ensemble, using the test data performance. Namely, the SuperLearner provides coefficients to inform of each learner’s weight, $A_{i}$, or importance in the overall ensemble. By default, the weights are always greater than or equal to 00 and sum to $1$. This approach is asymptotically as accurate as the best possible prediction algorithm tested. Our results show that the model produced by BGLM using Eq. A3 obtains the highest weight ($A_{1}=0.46$), followed by StepAIC ($A_{2}=0.25$), GLM using Eq. A1 ($A_{3}=0.16$), GAMs using Eq. A1 ($A_{4}=0.06$), Eq. A3 ($A_{5}=0.02$) and Eq. A2 ($A_{6}=0.01$). The rest of the models have a non-zero weight $>1\%$.
@@
These coefficients inform us on which one is the best model and reflects the weighted average of multiple models. The metric we use to define the goodness of our results is based on the minimization of the RMSE in the SuperLearner algorithm. To further measure the goodness of our results, along with the SuperLearner RMSE, we use the following metrics: the Pearson correlation coefficient, $r$, between $z_{obs}$ and $z_{pred}$, the mean squared error, MSE, and the bias.
@@
Our results show that the accuracy of the prediction is stable for the majority of partitions. However, we observe a higher RMSE (0.62) and a slightly smaller correlation coefficient of 0.91 for a small number of partitions. This is indeed natural due to the large heterogeneity of the data and the relatively small sample size. It should be noted that the default SuperLearner implementation of the methods mentioned above, namely GAM, GLM, and BGLM, utilizes a default formula for fitting the data, a smoothed linear combination of the explanatory variables. We have edited the definitions of these functions in R to ensure they use Eq. A1, A2, and A3 as the fitting functions.

### v2

@@ 4 Summary and Conclusion
We tested an ensemble with several algorithms to determine the redshift of GRBs based on characteristics related to their prompt and, for the first time, the plateau emission. Some afterglow characteristics (duration of the plateau emission, its flux, and the decay after the plateau) are among the most powerful predictors. A successful application of these methods to the Swift existing GRB catalog can nearly double the sample of GRBs with observed redshift and observed plateau, thus yielding a much larger sample. This, in turn, has allowed us to determine the LF and the density rate evolution of GRBs with unprecedented accuracy. Using a sample of 147 GRBs from the Swift satellite with known redshifts, we demonstrate the SuperLearner’s capabilities to provide accurate $z_{pred}$ that agree with the $z_{obs}$ with a Pearson correlation coefficient of $r=0.93$, and MSE=$0.21$. This result makes GRBs powerful distance estimators, improving upon other existing methods. Indeed, we obtained an increase of $61\%$ in the Pearson correlation of the $z_{pred}$ and $z_{pred}$. Instead, previous works, which omitted plateau parameters and used only one method, achieved a correlation of $r=0.57$. Results from applying this method to a large number of LGRBs have cast light on the controversy on the discrepancy between the density rate of LGRBs and the cosmological SFR and to accurately measure the GRB LFs. Our results mirror similar deviations in the formation rates of LGRBs found by Tsvetkova et al. (2021); Lloyd-Ronning et al. (2019); Yu et al. (2015); Petrosian et al. (2015); Pescalli et al. (2016). Notably, Pescalli et al. (2016) suggests a deviation of 2.6 sigma when considering the star formation history of Madau & Dickinson (2014). Furthermore, Dainotti et al. (2021) also arrived at a similar conclusion regarding SGRBs at low redshifts, even though the delayed SFR has been accounted for. Our primary conclusion, which aligns with findings reached by Petrosian & Dainotti (2023), is that the population of LGRBs at low redshift includes a component of compact merger progenitors, similar to the SGRBs. A different explanation for the excess of GRBs at low redshift, originally proposed by Yoon et al. (2006), is related to the increase in metallicity with cosmic time, which prevents a larger fraction of massive stars from producing GRBs.
@@ 5 Acknowledgement
We are particular grateful to Michal Ostrowski for his precious discussion during the data analysis and results, to Trevor Nelson for his preliminary analysis during the summer internship program at Stanford University supported by SLAC, to Simona Scimemi for the preliminary analysis related only to prompt emission variables during the winter internship 2018 at INAF Bologna, to Monica Brora on the very preliminary analysis on random forest. We also grateful to Macjei Bilicki for the useful discussion on the parameter space division. We thank particularly Douglas Figberg for his interesting discussion on selection biases related to this procedure and the way on how to solve them. M.G.D is particularly grateful for the support received as RIKEN pionieer research during her stay in Riken. M.G.D is grateful to the Gill Ferrandes and Donald Warren for their interesting discussions on the method of SuperLearner. We are also grateful to Trevor Hastie for useful suggestions to avoid overfitting and piece of the code for the 10-fold cross-validation. We are also grateful for the initial discussion on the project to Dmitry Maleshev. M.G.D. acknowledges the funding from MINIATURA2, Numer: 2018/02/X/ST9/03673. M.G.D also acknowledges Prof. Peter Behroozi for the insightful discussion on how to avoid overfitting in our statistical learning models. M.G.D and A.N also thank Nikita Khatiya, Dhruv Bal and Dr. Dieter Hartmann for extremely valuable feedback on the DRE analysis. This research was also support by the Visibility and Mobility module of the Jagiellonian University (Grant number: WSPR.WSDNSP.2.5.2022.5) and the NAWA STER Mobility Grant (Number: PPI/STE/2020/1/00029/U/00001). We acknowledge funding from Swift GI Cycle XIX, grant number 22-SWIFT22-0032. This work was also supported by the Polish National Science Centre grant UMO-2018/30/M/ST9/00757 and by Polish Ministry of Science and Higher Education grant DIR/WK/2018/12.
@@ A.1 Statistical learning methods
We have used predictive data mining, also called supervised learning. It is based on prior knowledge of a ‘training’ data set on which we can build models to predict the relations in a new “test” data set. Examples are regression tools, such decision trees, random forests, and gradient boosting. These allow us to estimate complicated non-linear relationships between the response and predictors as well as high-order interactions between predictors. Such non-parametric methods are very powerful when the data set contains many observations. However, they suffer from the so-called “curse of dimensionality”, which limits the number of predictors one can efficiently estimate for a small sample size. As a result, the fully non-parametric machine learning methods allow using only a limited number of predictors, e.g., when estimating redshifts based on small GRB training sets. On the other hand, when important predictors are identified by other methods or expert knowledge, these methods allow the construction of efficient predictors without assumptions on the underlying form of the relevant correlations among the variables. In our analysis, we have used supervised statistical learning tools that employ regression. Within regression methods, we have used parametric, semi-parametric, and non-parametric approaches. We give a brief description of each model below.
A.1.1 Predictors selection
@@
The GRB sample contains a total of nine predictors. Four belong to the prompt: $Fluence$, $T_{90}$, $\log(Peakflux)$, and $Photon$ $Index$. The rest belong to the plateau and afterglow: $T_{a}$, $F_{a}$, $\alpha$, $\beta$, and $NH$. To pick out the best predictors of redshift among these 9 variables, we applied the Least Absolute Shrinkage and Selection Operator (LASSO) predictor selection technique. The LASSO technique applies a shrinkage method for linear regression by forcing the $\ell^{1}$ norm (sum of the magnitude of all vectors in the given space) of the result vector to be $\leq\lambda$, where $\lambda$ is a positive number known as the tuning parameter. This form of penalization enables the model to choose specific predictors while nullifying others by assigning their coefficients a value of 0 (Tibshirani, 1996). The tuning parameter determines the degree of shrinkage applied to the estimated vector. Consequently, when using fewer predictors, the model becomes more interpretable and generally exhibits lower prediction errors than the full model. For our analysis, we use the GLMNET function with the LASSO selection predictors (Hastie et al., 2017; Tibshirani et al., 2012). We pick the $\lambda.1se$ value, the maximum $\lambda$ value for which the error is within one standard deviation (Friedman et al., 2010a), and its corresponding coefficients for the predictors s. Thus, we drop $T_{90}$ and $Fluence$ while keeping the rest of the 7 variables. The coefficients assigned by LASSO to each of them are displayed in Fig. 6, and we choose only the non-zero coefficient predictors. The LASSO selected predictors are also presented in Fig. 1.
@@
Figure 6: LASSO-based predictors selection, showing the weights assigned to the different variables.
@@ A.1.2 The Generalized Additive Model
In the semi-parametric generalized additive models (GAM), $Y$ is related to the predictors, $x_{1},\ldots,x_{p}$ through the additive model $f(x_{1},\ldots,x_{p})=\beta_{0}+\sum_{i=1}^{p}{\beta_{i}f_{i}(x_{i})}$. The functions $f$ may have some specified parametric form (e.g., a polynomial) or be estimated non--parametrically, simply as ‘smooth functions.’ In this method, interactions between different predictors can be included. An interaction may arise when there are two or more predictors, and the influence of some of them on the response variable depends on the other predictors. In practice, interactions make it more difficult to predict the consequences of changing the value of a variable. If any smooth functions were allowed in model fitting, then the maximum likelihood estimation of such models would result in over-fitting. Thus, the models are usually fit by penalized likelihood maximization, in which the model is modified by adding a penalized likelihood estimation, namely a penalty for each smooth function, penalizing its irregularities.
We use GAM to determine an analytical formula that best estimates the $\log(z+1)$ value. It should be noted that we present results in log scale because of the log scale transformations we applied to the majority of the variables and because the best model is indeed trained in $\log(1+z)$. In addition, Z = $z+1$ is a more natural parametrization of the cosmological evolution. The parameters of this formula are the LASSO selected predictors and their cross multiplicative terms (terms defined as the products among the 7 chosen variables). A total of 16,511 formulas were generated considering all possible combinations of all seven predictors and the product of the pairs of these variables.
@@
{dmath}
log(z+1) = (log(Peak)^2 + log(NH)^2 + PI^2 + β^2 + log(Peak) + α)^2 + log(F_a) + log(T_a) + β+ PI + log(NH) + log(F_a)^2 + log(T_a)^2 + α^2
{dmath}
log(z+1) = (log(Peak)^2 + log(NH)^2 + β^2 + log(Peak) + α)^2 + log(F_a) + log(T_a) + β+ PI + log(NH) + log(F_a)^2 + log(T_a)^2 + α^2 + PI^2
{dmath}
log(z+1) = (log(Peak)^2 + log(NH)^2 + β^2 + log(Peak) + log(NH))^2 + log(F_a) + log(T_a) + α+ β+ PI + log(F_a)^2 + log(T_a)^2 + α^2 + PI^2
@@ A.1.3 Step-AIC
Step-AIC is a parametric approach that chooses algorithms based on the Akaike Information Criterion (AIC) Sakamoto et al. (1986); Venables et al. (2002), which, given a collection of models, estimates the quality of each model relative to each other. The StepAIC model works by fitting a formula to which it adds (or subtracts) predictors iteratively while minimizing the AIC.
@@ A.1.4 The Generalized Linear Model
Generalized Linear Model (GLM) Nelder & Wedderburn (1972) is a parametric method that estimates the mean of the response variable’s distribution using a linear model of the predictors. The response variable is connected to the desired target variable (GRB redshift) via a link function. This is an extension of the ordinary linear model. Using the appropriate link functions, we can consider non-Gaussian target variable distributions, such as Poisson or Gamma functions. GLM adopts the iteratively weighted least squares method for maximum likelihood estimation to fit these functions. Here, we assign all three formulas (Eq. A.1.2, A.1.2, and A.1.2) as the desired fitting functions, while the link function is Gaussian.
@@ A.2 Outlier removal with M-estimator
To remove outliers that might not represent the full sample, we apply the Robust Linear regression Method (RLM) with the M-estimator. This is an alternative technique to the ordinary least squares method, which for a univariate response variable, $Y$ and a vector $X=(X_{1},\ldots,X_{p})$ of predictors $X$, attempts to fit the following relation: $f(x_{1},\ldots,x_{p})=\beta_{0}+\sum_{i=1}^{p}{\beta_{i}x_{i}}$. The ordinary least squares method attempts to minimize the square of the residuals (called the $L_{2}$ norm regression). However, this gives outliers of the data set a higher weight, which significantly affects the results of the regression fit. In contrast, the M-estimator attempts to minimize the sum of a function of the residuals. The function chosen for our analysis is the Huber Function (Huber, 1964).
@@
Iterated re-weighted least squares do the M-estimator fitting, and the weights, ranging from 0 to 1, are assigned based on the residuals (Huber, 1996). RLM is used for the detection of highly influential observations. We are using the implementation of RLM as described in the MASS package of R (Venables & Ripley, 2002). We call rlm() function and specify the Eq. A.1.2 based on which the fitting is performed. To ensure the rlm() function uses the M-estimator method for assigning weights, we specify the parameter method=‘M’, where M indicates the M-estimator techniques. For further discussion on robust statistical methods see Stromberg (2004); Hampel et al. (2011); Marazzi (1993).
@@
Here, choose a cut-off weight of 0.5, eliminating GRBs with a weight $\leq 0.5$. The threshold of 0.5 was chosen to ensure we remove less than 5% of our sample as outliers. With the M-estimator, the sample size is reduced from 150 to 147.
A.3 The SuperLearner and the Relative importance of the predictors
@@
SuperLearner is an algorithm that uses internal 10-fold cross-validation (10fCv) to estimate the performance of multiple statistical learning models or the same model with different settings (Van der Laan et al., 2007; Naimi & Balzer, 2018). It minimizes the estimated risk (a measure of model accuracy) by minimizing the RMSE. To test the performance of our results and estimate the prediction error, besides the 10fCV, we also used an external layer of cross-validation called nested cross-validation. Namely, we performed 10fCV repeatedly $100$ times, with the folds of the cross-validation step being randomized for each iteration. This type of cross-validation involves partitioning the sample into complementary subsets (in this case, $10$), fitting the model on 9 out of $10$ folds (training sets) of the full data set, and validating the analysis on the remaining 1 testing set. The procedure is repeated iteratively so that each set of the $10$ will be used as a test set and the remaining as training sets, and the prediction results are averaged over the number of runs (in this case $100$). This step allows us to remove the randomness inherent to the cross-validation procedure and stabilize the results. Then, SuperLearner creates an optimal weighted average of those models, e.g., an ensemble, using the test data performance. Namely, the SuperLearner provides coefficients to inform of each learner’s weight, $A_{i}$, or importance in the overall ensemble. By default, the weights are always greater than or equal to 00 and sum to $1$. This approach is asymptotically as accurate as the best possible prediction algorithm tested. Our results show that the model produced by BGLM using Eq. A.1.2 obtains the highest weight ($A_{1}=0.46$), followed by StepAIC ($A_{2}=0.25$), GLM using Eq. A.1.2 ($A_{3}=0.16$), GAMs using Eq. A.1.2 ($A_{4}=0.06$), Eq. A.1.2 ($A_{5}=0.02$) and Eq. A.1.2 ($A_{6}=0.01$). The rest of the models have a non-zero weight $>1\%$.
@@
These coefficients inform us which is the best model and reflects the weighted average of multiple models. The metric we use to define the goodness of our results is based on the minimization of the RMSE in the SuperLearner algorithm. To further measure the goodness of our results, along with the SuperLearner RMSE, we use the following metrics: the Pearson correlation coefficient, $r$, between $z_{obs}$ and $z_{pred}$, the mean squared error, MSE, and the bias.
@@
Our results show that the accuracy of the prediction is stable for the majority of partitions. However, we observe a higher RMSE (0.62) and a slightly smaller correlation coefficient of 0.91 for a small number of partitions. This is indeed natural due to the large heterogeneity of the data and the relatively small sample size. It should be noted that the default SuperLearner implementation of the methods mentioned above, namely GAM, GLM, and BGLM, utilizes a default formula for fitting the data, a smoothed linear combination of the predictors. We have edited the definitions of these functions in R to ensure they use Eq. A.1.2, A.1.2, and A.1.2 as the fitting functions.